### v4

# SemRel2024: A Collection of Semantic Textual Relatedness Datasets for 13 Languages

Abstract
Exploring and quantifying semantic relatedness is central to representing language and holds significant implications across various NLP tasks. While earlier NLP research primarily focused on semantic similarity, often within the English language context, we instead investigate the broader phenomenon of semantic relatedness. In this paper, we present SemRel, a new semantic relatedness dataset collection annotated by native speakers across 13 languages: Afrikaans, Algerian Arabic, Amharic, English, Hausa, Hindi, Indonesian, Kinyarwanda, Marathi, Moroccan Arabic, Modern Standard Arabic, Spanish, and Telugu. These languages originate from five distinct language families and are predominantly spoken in Africa and Asia – regions characterised by a relatively limited availability of NLP resources. Each instance in the SemRel datasets is a sentence pair associated with a score that represents the degree of semantic textual relatedness between the two sentences. The scores are obtained using a comparative annotation framework. We describe the data collection and annotation processes, challenges when building the datasets, baseline experiments, and their impact and utility in NLP.
1 Introduction
Figure 1: SemRel2024 languages and language families.
Characterising the relationship between two units of text is an important component of constructing text representations. Within this context, semantic textual relatedness (STR) aims to capture the degree to which two linguistic units (e.g., words or sentences, etc.) are close in meaning Mohammad (2008); Mohammad and Hirst (2012). Two units may be related in a variety of different ways (e.g., by expressing the same view, originating from the same time period, elaborating on each other, etc.). On the other hand, semantic textual similarity (STS) considers only a narrow view of the relationship that may exist between texts (such as equivalence or paraphrase) which does not incorporate other dimensions of relatedness such as entailment, topic or view similarity, or temporal relations Abdalla et al. (2023); Agirre et al. (2013a). For example, ‘I caught a cold.’ and ‘I hope you feel better soon.’ would receive a low similarity score, despite the two being very related. In this work, we investigate the broader concept of semantic textual relatedness.
STR is central to understanding meaning in text Hasan and Halliday (1976); Miller and Charles (1991); Morris and Hirst (1991) and its automation can benefit various downstream tasks such as evaluating sentence representation methods, question answering, and summarisation Abdalla et al. (2023); Wang et al. (2022).
Table 1: Examples of sentence pairs and their corresponding scores (from 0 to 1) in the various SemRel2024 languages. Examples are sorted by score and rows with higher degrees of relatedness are lighter colored. The translations can be found in the Appendix.
Prior NLP work has mainly focused on textual similarity, largely due to a dearth of relatedness datasets. Of the existing STR and STS datasets, most are in the English. The few STR and STS resources which exist for non-high resource languages are composed of word-level or phrase-level pairings. In this work, we curate 13 new monolingual STR datasets [^1] Our team also created data for Punjabi. However, the sentence-pair selection procedure for it was markedly different than what we used for other languages. Therefore we do not include it here. for Afrikaans (afr), Amharic (amh), Modern Standard Arabic (arb), Algerian Arabic (arq), Moroccan Arabic (ary), English (eng), Spanish (esp), Hausa (hau), Hindi (hin), Indonesian (ind), Kinyarwanda (kin), Marathi (mar), and Telugu (tel).
The datasets are composed of sentence pairs, each assigned a relatedness score between 0 (completely unrelated) and 1 (maximally related). With the aim of curating diverse STR datasets, the pairs of sentences were first selected from pre-existing datasets covering various topics and formality levels, e.g., news data, Wikipedia, and conversational data. Additionally, we selected pairs with a large range of expected relatedness values by considering lexical overlap, contiguity, topic coverage, and random pairings. To generate the relatedness scores, the sentence pairs were then annotated by native speakers who performed comparisons between different pairs of sentences using Best-Worst Scaling (BWS) Louviere and Woodworth (1991). BWS is known to avoid common limitations of traditional rating scale annotation methods Kiritchenko and Mohammad (2016, 2017). The annotation process led to the high reliability of the final relatedness rankings in the different SemRel datasets. Our main contributions are as follows:
  1. 1.
    We present the first benchmark on semantic distance (similarity or relatedness) that includes low-resource African and Asian languages from five different language families (see Figure 1). Although Africa and Asia are home to over 5,000 languages from over 20 language families and have the highest linguistic diversity, there is little publicly available data on these languages.
  2. 2.
    We discuss general and language-specific challenges related to the data collection and annotation of the SemRel datasets.
  3. 3.
    We present baseline experiments conducted in different monolingual and crosslingual settings to demonstrate the usefulness and potential of our dataset collection.
Lang. Curation technique Data Sources
afr Overlap, Random selection, Manual check News data, reviews, recipes, blogs.
amh Overlap, Similarity, Random selection, Manual check News data, crawling.
arb Overlap, Contiguity, Random selection, Manual check Ted talk subtitles, news data.
arq Overlap, Contiguity, Random selection, Manual check YouTube comments, conversational data.
ary Overlap News data.
eng Overlap, Similarity, Paraphrases, Contiguity, Randomness Book reviews, news data, tweets, other.
esp Overlap, Contiguity, Similarity Movie reviews, news data, other.
hau Overlap News data.
hin Overlap, Similarity, Contiguity, Paraphrase, Randomness News data, other.
ind Overlap Wikipedia, news data.
kin Overlap News data.
mar Overlap, Similarity, Contiguity, Paraphrase, Randomness News data, other.
tel Overlap, Similarity, Contiguity, Paraphrase, Randomness News data, other.
Table 2: The curation techniques used for data creation. We list the main textual sources present in the datasets we used for instance creation. More details are shared in Section 3.1.2.
To promote research in the field of semantic relatedness, we publicly released the SemRel2024 datasets as part of a shared task that attracted a large number of participants interested in low-resource languages. [^2] See https://semantic-textual-relatedness.github.io for more details.
2 Related Work
The field of semantic relatedness in natural language processing covers a variety of approaches and techniques designed to measure the closeness in meaning between units, specifically words Miller (1994), or sentences Abdalla et al. (2023).
Most prior work focuses on STS, a narrower subset of STR, and often only covers high-resource languages such as English Agirre et al. (2012, 2013b, 2014, 2015, 2016); Marelli et al. (2014), Arabic, German, Spanish, Turkish Ahmed et al. (2020); Cer et al. (2017b), and Italian Glavaš et al. (2018) with the only exception being Finnish, Slovene, Croatian Glavaš et al. (2018); Armendariz et al. (2020) and Farsi Vulić et al. (2020). To overcome the scarcity of available resources, Tang et al. (2018) proposed sentence-level, encoder-based methods leveraging English data to create Arabic, Spanish, Thai, and Indonesian datasets, whereas Pandit et al. (2019) use traditional data augmentation methods to create Bangla data.
By comparison, this work is focused on the creation of resources for sentence-level STR in multiple low-resource languages. Here, the few works which exist for non-high-resource languages are at the word level (e.g., Yum et al., 2021 for Korean). To our knowledge, the only corpora specially designed for semantic textual relatedness between pairs of sentences was created by Abdalla et al. (2023) for English. Abdalla et al. (2023) curated a dataset of 5,500 English sentence pairs annotated using a comparative annotation framework. Their dataset has since been used to evaluate embedding approaches Wang and Li (2022) and other methods Wang et al. (2022). The core of Abdalla et al. (2023) approach serves as the model for data annotations in this project. However, our work additionally explores new ways of data collection–curation, and several challenges had to be addressed when working with less-resourced languages.
3 STR Data
3.1 Data collection
A key step in the data creation process was identifying sources of text for each language and selecting sentence pairs. This was particularly challenging for low-resource languages such as Hausa, Kinyarwanda, and Algerian Arabic. Since arbitrarily selecting sentences and pairing them would lead to many unrelated instances, we relied on several heuristics, discussed in Section 3.1.1, to ensure a wide range of scores for each language. Since these methods are highly corpus- and language-specific, the approaches used per language were determined by native speakers. We provide the data origin and the pairing approaches used for each language in Section 3.1.2. The composition of the resulting dataset is summarised in Table 6 and the distribution of the relatedness scores across the datasets are illustrated in Figure 2.
3.1.1 Sentence pairing heuristics
Given a set of texts in a target language, careful consideration was given to the construction of sentence pairs to ensure that the pairs would exhibit relatedness scores varying from completely unrelated to very related. Since random selection would result in many unrelated pairs, we paired sentences mainly based on five methods previously defined by Abdalla et al. (2023) (described below). In cases where the pairs produced by the five methods were qualitatively judged to be insufficiently varied, we manually selected some instances to balance the data so that we have sufficient number of instances for each band of relatedness (high, medium, low, or unrelated).
Lexical overlap
Pairs are selected with various amounts of lexical overlap. That is, one or more words/tokens in common, with or without using TF-IDF normalisation. This method is expected to produce a wide range of relatedness values, and was used in most low-resource languages.
Contiguity/Entailment
We select pairs of sentences that appear one after the other in a paragraph or a social media thread. This method is likely to produce pairs of sentences that are somewhat related and can contribute to representing the low to medium ranges of relatedness.
Paraphrases or Machine Translation (MT) paraphrases
This method consists of selecting pairs of sentences from paraphrase or MT data. For MT, we pivot across the translation and back to the source language to generate a new sentence and pair it with the original. However, many low-resource Asian and African languages lack reliable MT resources.
Semantically similar instances
Semantically similar sentences are selected from a publicly available dataset such as the STS dataset by Cer et al. (2017a) or manually identified by a native speaker in order to include highly related instances and balance the dataset.
Random selection
Random sentences are selected. This method is expected to represent the low to medium ranges of relatedness.
Manual check
In cases where the pairs produced by the above methods were qualitatively judged to be insufficiently varied, instances were manually selected to balance the data so that there were sufficient number of instances for each band of relatedness (high, medium, low, or unrelated). This can apply to any range of relatedness (i.e., high, medium, low, or unrelated).
Language afr amh arb arq ary eng esp hau hin ind kin mar tel
#Ann/tuple 2 4 2-3 2 2 2-4 2-4 2-4 4 2 2 2-3 4
SHR train/dev 0.85 0.90 0.86 0.64 0.77 0.84 0.70 0.74 0.93 0.68 0.74 0.92 0.79
SHR test 0.85 0.90 0.86 0.64 0.77 0.80 0.70 0.74 0.94 0.68 0.74 0.96 0.96
Table 3: SHR (split-half reliability) scores for each of the dataset splits and numbers of unique annotations per tuple (#Ann/tuple). As some languages (eng, hin, mar, and tel) had splits annotated in separate annotation efforts (instead of one combined annotation), we report the SHR scores for both.
3.1.2 Data curation
Since most of the SemRel languages are low-resource, the domain, (in)formality, and diversity of the sentence pairs were highly dependent on the publicly available corpora. We aimed to collect datasets with average-length sentences, free of offensive utterances, and as diverse as possible. As such, data instances were extracted for each language using a tailored combination of the heuristics described in Section 3.1.1. We used further pre-processing, post-processing, and data analysis methods (discussed below) to avoid incoherence and unnaturalness.
English and Spanish
As English and Spanish are high-resource languages, we sampled sentences from various sources to capture a wide variety of sentence structure, formality, and grammaticality in texts. As shown in Table 6, we paired sentences in a number of ways that include lexical overlap, entailment, similarity and paraphrases. The English dataset includes sentences that have the same meaning but a different formality collected from the Formality dataset Rao and Tetreault (2018), tweets Mohammad et al. (2017), paraphrases from machine translation systems extracted from the ParaNMT dataset Wieting and Gimpel (2017), book reviews from Goodreads Wan and McAuley (2018), pairs of premises and hypotheses from the SNLI dataset Bowman et al. (2015), and semantically similar sentences Cer et al. (2017a).
Similarly, we selected Spanish sentence pairs from semantic similarity datasets Agirre et al. (2014, 2015); Cer et al. (2017a), entailment datasets such as SICK-es Huertas-Tato et al. (2021) and NLI-es Araujo et al. (2022). Additionally, we sampled contiguous sentences from the summarisation dataset XL-Sum Hasan et al. (2021), DiscoEval Spanish Araujo et al. (2022), and questions of different types from Spanish QC Á. García Cumbreras et al. (2006).
Arabic Variations: Modern Standard Arabic, Algerian, and Moroccan Arabic
Arabic is known for diglossia Ferguson (1959), meaning that Arabic varieties are used for different contexts. For instance, Modern Standard Arabic is usually used in formal and academic communication while dialects are typical for conversational settings. The various sources of the Arabic data are somewhat reflective of the distinct language usage scenarios.
Therefore, for Modern Standard Arabic (MSA), we used two datasets from different domains: TED Talk subtitles Zong (2015) on science, society, and art and news articles on economics Al-Dulaimi (2022). In addition to sentences with lexical overlap, we selected contiguous sentences in Ted Talk subtitles to include different degrees of relatedness, and as some sentences in the subtitles were slightly ungrammatical, we corrected them based on the standard Arabic grammar rules. For Algerian Arabic, we used CalYou Abidi et al. (2017), a dataset composed of YouTube comments collected from major Algerian YouTube channels by 2017, and the Algerian instances spoken in two major Algerian towns (Algiers and Annaba) present in PADIC: Parallel Arabic Dialect Corpus Meftouh et al. (2015). We used lexical overlap to pair sentences, picked contiguous ones in a conversation in PADIC, and added randomly or manually selected sentence pairs to balance the relatedness score distribution in the dataset. For both MSA and Algerian Arabic, we allowed short sentences as Arabic is highly inflectional. For Moroccan Arabic, we used headlines from the Goud.ma dataset introduced by Issam and Mrini (2022) and the Moroccan Arabic sentences were paired based on lexical overlap.
Afrikaans
The Oscar dataset Ortiz Suárez et al. (2020) was used as basis for the Afrikaans corpus. We chose sentences from news articles, blogs, reviews, and recipes. We also excluded sentences from religious texts and academic articles after observing that these did not produce high-quality pairs. We further excluded a number of advertorial texts that appear to be low-quality translations. All instances were then manually assessed for grammar and ungrammatical sentences were discarded. Sentences were paired if they had an overlap of at least five tokens and at least three non-overlapping tokens with matches within the same article only. Random sentence pairs were also included to calibrate the dataset.
Amharic, Hausa, Kinyarwanda
For Amharic, we paired sentences present in news articles from different Ethiopian news outlets Yimam et al. (2021). Similarly, the Hausa and Kinyarwanda datasets include pairs of sentences from news articles collected by Abdulmumin and Galadanci (2019) and Niyongabo et al. (2020), respectively. Sentences shorter than five words and longer than 20 were excluded and pairs were created using lexical overlap. Additionally, for Amharic, we excluded sentences with mixed languages to avoid confusing the annotators.
Indonesian
For Indonesian, we collected sentences from Wikipedia texts present in the ROOTS split Laurençon et al. (2022); Setya and Mahendra (2018) and the IndoSum Kurniawan and Louvan (2018) datasets. IndoSum is a human-written summarization dataset consisting of pairs of news articles with their abstractive summaries. We parsed both corpora at a sentence-level and only selected sentences that are composed of five to fifteen words.
Hindi, Marathi, and Telugu
As these languages lack publicly available resources, especially Marathi and Telugu, we used the Mukhyansh dataset only Madasu et al. (2023) to create the sentence pairs. It is composed of news headlines and their corresponding articles and is diverse in nature. We created instances using lexical overlap, paraphrase generation, contiguous sentence selection, and random sentence selection to balance the data.
Figure 2: Violin plots representing the distributions of the relatedness scores. For instance, the distribution of the Arabic (arb) dataset is unimodal, Marathi’s (mar) is bimodal, and the Indonesian (ind) dataset’s is trimodal.
3.2 Data annotation and challenges
Data afr amh arb arq ary eng esp hau hin ind kin mar tel
Train - 992 - 1,261 924 5,500 1,562 1,736 - - 778 1,200 1,170
Test 375 171 595 583 426 2,600 600 603 968 360 222 298 297
Dev 375 95 32 97 71 250 140 212 288 144 102 293 130
Total 700 1,258 627 1,941 1,421 8,350 2,302 2,551 1,256 504 1,102 1,791 1,597
Table 4: Number of instances in the training, dev, and test sets for the different datasets. The languages with no training data (afr, arb, hin, ind) were only used in unsupervised and cross-lingual settings.
Annotation process
Similarly to Abdalla et al. (2023), we used BWS to annotate our data instances and generate an ordinal ranking of instances [^3] The tuples were generated using the BWS scripts provided by Kiritchenko and Mohammad (2017): http://saifmohammad.com/WebPages/BestWorst.html.. Although pairwise comparisons are more reliable than simply labelling the sentence pairs as related or unrelated, it is a time-consuming process if performed on a large dataset as it requires $N\times N=N^{2}$ comparisons if performed on a dataset of $N$ instances. Best-worst scaling mitigates this issue according to Kiritchenko and Mohammad (2017) as it leads to reliable scores from about $2\times N$ comparisons of 4-instance tuples.
BWS requires fewer labels Louviere and Woodworth (1991), in our case, given four instances (i.e., pairs of sentences) $p_{i}$ with $40\leq i<$ 4, for a tuple: $\langle p_{0},p_{1},p_{2},p_{3}\rangle$, if $p_{0}$ is marked as most related and $p_{3}$ as least related, then we know that $p_{0}>p_{1}$, $p_{0}>p_{2}$, $p_{0}>p_{3}$, $p_{1}>p_{3}$, and $p_{2}>p_{3}$ ($<$ and $>$ refer to less and more related, respectively). We then use these inequalities to compute real-valued scores that consist of the fraction of times a pair $p_{i}$ was chosen as the most related minus the fraction of times $p_{i}$ was chosen as the least related. Then, an ordinal ranking of sentence pairs is generated Orme (2009); Flynn and Marley (2014).
Furthermore, the notions of related and unrelated have fuzzy boundaries with no singular accepted definition in the literature. Different people and different language cultures may have several intuitions of where such a boundary exists. Therefore, by using comparative annotations and relying on the intuitions of fluent speakers for each language to choose between sentence pairs, we can avoid ill-defined categories. This is in line with our goal of capturing common perceptions of semantic relatedness (i.e., what is believed by the vast majority) instead of “correct” or “right” rankings.
Instructions
We selected native speakers to annotate the sentence pairs. Then, given a set of four sentence pairs, annotators were tasked with reporting on their relative relatedness. Concretely, given 4 sentence pairs, each of the form [sentence A, sentence B], the task was to select the sentence pair that is the most related (i.e., sentence A is closest in meaning to sentence B) and the sentence pair that is the least related (i.e., sentence A is farthest in meaning to sentence B). The full instructions can be found in the Appendix.
In the guidelines, it was noted that sentence pairs that are more specific in what they share tend to be more related than sentence pairs that are only loosely about the same topic. Furthermore, if one or both sentences have more than one interpretation, the annotators have to consider the closest meanings.
Overall, by manually examining the annotations, we noted that the BWS framework does lead to more robust annotations. However, a downside is the fact that annotating one instance could take more than one minute and the task can be challenging since many instances to be compared can be similarly (un)related.
Annotation reliability
In Table 3, we report the number of annotators and the split-half reliability (SHR) Cronbach (1951); Kuder and Richardson (1937) scores for each of the datasets. SHR measures the degree to which repeating the annotations results in similar relative rankings of the instances. First, it splits the 4-tuple annotations into two bins. Then, the annotations for each bin are used to generate two different independent relatedness scores, and the Spearman correlation between the two sets of scores is calculated to estimate the closeness of the two rankings. A high correlation indicates that the annotations are reliable. This process is repeated 1,000 times and the correlation scores are averaged similarly to Abdalla et al. (2023). Overall the scores in Table 3 vary between 0.64 and 0.96, which indicates a high annotation reliability.
Disagreements
We inspected annotators with large disagreements to ensure the annotation procedure was correctly followed (i.e., their annotations made sense for native speakers). Very strong disagreements would serve as a red flag of poor data quality resulting in a more thorough review of the annotator quality. Hence, as a sanity check, we looked to see that sentences with high relatedness scores were more semantically related than those with low relatedness scores. The specific procedure for ensuring data quality depends on the annotation procedure of the team (e.g., those using AMT vs those who did not). Note that disagreements were not deleted, as they can serve as a useful signal for BWS. That is, for a tuple $\langle p_{0},p_{1},p_{2},p_{3}\rangle$, when annotators disagree on what is most related (e.g., $p_{1}$ or $p_{3}$), then it is an indication that $p_{1}$ and $p_{3}$ may be semantically close to each other. As all tuple annotations (twice the number of instances) are used to determine the final scores of the sentence pairs, this disagreement would lead to the two pairs ($p_{1}$ and $p_{3}$) getting scores that are close to each other. On the other hand, if a sentence pair consistently occurs in 4-tuples that have very low annotator agreement, then it is likely that the sentence pair is the source of disagreement. This can be due to various reasons such as the language use, code-switching, or the annotator’s familiarity with the topics discussed.
Besides sharing our datasets with the community, we also make the full 4-tuple annotations public.
3.3 Postprocessing and data quality control
Quality control
For our final dataset, we carried a data post-processing step to ensure that:
  • no instances are repeated;
  • the data does not include invisible characters, incorrectly rendered emoticons, or garbled encoding characters;
  • texts are fully anonymised (deleting emails and IDs if they occur, replacing @mentions with @<username>, and replacing any URLs with non-identifiable placeholders);
  • the data does not include a high amount of expletives or inappropriate language; and
  • the data is balanced.
Manual Spotchecks
Finally, a team of native speakers manually spotchecked the scores to make sure that the relatedness scores made sense and to supplement the quantitative evaluation based on SHR.
4 Experiments
4.1 Data
We use the splits reported in Table 4. For the languages without training data (afr, arb, hin, ind), we only report experiments in unsupervised and crosslingual settings. For English, we use the STR-2022 dataset Abdalla et al. (2023) for training and we use our newly created dataset for testing.
4.2 Experimental setup
afr amh arb arq ary eng esp hau hin ind kin mar tel
Overlap 0.71 0.63 0.32 0.40 0.63 0.67 0.67 0.31 0.53 0.55 0.33 0.62 0.70
Unsupervised (Multilingual)
mBERT 0.74 0.13 0.42 0.37 0.27 0.68 0.66 0.16 0.62 0.50 0.12 0.65 0.66
XLMR 0.56 0.57 0.32 0.25 0.17 0.60 0.69 0.04 0.51 0.47 0.13 0.60 0.58
Unsupervised (Monolingual)
AfroXLMR 0.45 0.40 0.18 - - 0.30 - 0.07 - - 0.16 - -
ALBETO - - - - - - 0.62 - - - - - -
AmRoBERTa - 0.72 - - - - - - - - - - -
ARBERT - - 0.56 - - - - - - - - - -
arb BERT - - 0.31 - - - - - - - - - -
BETO - - - - - - 0.68 - - - - - -
DziriBERT - - - 0.43 - - - - - - - -
Indic-BERT - - - - - - - - 0.40 - - 0.41
MARBERT - - 0.29 - - - - - - - - - -
RoBERTa-BNE - - - - - - 0.66 - - - - - -
HauRoBERTa - - - - - - - 0.12 - - - - -
Supervised
LaBSE - 0.85 - 0.60 0.77 0.83 0.70 0.69 - - 0.72 0.88 0.82
Crosslingual
LaBSE 0.79 0.84 0.61 0.46 0.40 0.80 0.62 0.62 0.76 0.47 0.57 0.84 0.82
Table 5: Spearman correlation scores for different fine-tuned models in the three settings that we describe (supervised, unsupervised, and crosslingual) in addition to a simple lexical overlap baseline (Overlap).
We report the Spearman correlation scores between the predicted labels and the gold standard ones for the different languages in three main settings:
  • Supervised systems trained on the labeled training datasets provided.
  • Unsupervised systems developed without the use of labeled datasets pertaining to semantic relatedness or semantic similarity between units of text of more than two words long in any language.
  • Crosslingual systems developed without the use of any labeled semantic similarity or semantic relatedness datasets in the target language and with the use of data from at least one other language included in SemRel.
In our experiments, we use:
  • a simple baseline based on the number of shared words (lexical overlap),
  • sentence embeddings (LaBSE Feng et al. (2020), SentenceBERT Reimers and Gurevych (2019)), and
  • standard encoder-based embeddings.
4.3 Lexical Overlap
As shown in Table 5, we report a simple lexical overlap baseline which consists of the Dice coefficient between two sentences A and B: the number of unique unigrams occurring in both sentences, adjusted by their lengths Abdalla et al. (2023):
$\frac{2\times|unigram(A)\cap unigram(B)|}{|unigram(A)+unigram(B)|}$ (1)
4.4 Supervised and Crosslingual settings
We use LaBSE (Label Agnostic BERT Sentence Embeddings) Feng et al. (2020) which can map 109 languages into a shared vector space. With the embeddings covering all the SemRel languages, we report baseline results using the default hyperparameters set in the sentence-transformers repository [^4] https://github.com/UKPLab/used-transformers. Our experiments are conducted:
  • using the predefined setup without further fine-tuning,
  • by fine-tuning the LaBSE model on our training data using a cosine similarity loss.
We report the scores on the test sets in both setups in Appendix A, Table 12.
For the crosslingual baselines, we fine-tune LaBSE on the English training set and test on all the other datasets except English. On the other hand, when testing on the English dataset, we use the Spanish training set to fine-tune LaBSE.
4.5 Unsupervised settings
We used the standard encoder-based monolingual and multilingual language models on our datasets [^5] We use the standard models from HuggingFace Wolf et al. (2020).. We experiment with:
  • multilingual BERT (mBERT) Devlin et al. (2019), XLMRoberta (XLMR) Conneau et al. (2020) for all 13 languages,
  • monolingual models:
    • AfroXLMR Alabi et al. (2022) for Afrikaans, Amharic, Hausa, Kinyarwanda,
    • Indic-BERT Kakwani et al. (2020) for Hindi, Marathi, and Telugu,
    • BERT Devlin et al. (2019) for English,
    • MARBERT, ARBERT Abdul-Mageed et al. (2021) and Arabic BERT Safaya et al. (2020) for Arabic,
    • BETO Cañete et al. (2020), ALBETO Cañete et al. (2022), and RoBERTa-BNE Fandiño et al. (2022) for Spanish,
    • Amharic RoBERTa (AmRoBERTa) Yimam et al. (2021) for Amharic,
    • DziriBERT Abdaoui et al. (2021) for Algerian Arabic,
    • RoBERTa based model (HauRoBERTa) for Hausa Adelani et al. (2022).
We report the Spearman correlation scores with cosine similarity scores for all the BERT-based models in Table 5. Additional results using BERTScore Zhang* et al. (2020) for mBERT and XLMR are shared in the Appendix (see Table 13).
4.6 Experimental results
Table 5 shows the Spearman correlation scores for the three setups: supervised, unsupervised, and crosslingual for all thirteen languages. For the unsupervised models, we report the results using all pretrained models including mBERT Devlin et al. (2019) and XLMR Conneau et al. (2020). Additionally, we report the Spearman correlation for experiments with monolingual language-specific models for each language – models that have been trained in these specific languages.
For the general setup, we note that, except for Amharic and Kinyarwanda, mBERT outperforms XLMR in all languages by a significant margin. For Amharic, mBERT’s correlation score with the gold labels is 0.13, whereas XLMR’s is three times higher with a 0.57 correlation score. Surprisingly, even though Arabic is a high-resource language, the Spearman correlation score is relatively low in comparison to all the high-resource languages, with Spanish achieving the best results. This could be due to the size of the Arabic data being smaller.
For the language-specific models, the results are highly tied to the language. In cases such as Amharic for example, AmRoBERTa significantly improves the score by 0.27 points, whereas AfroXLMR hurts the performance for all African languages.
Similarly to the unsupervised setup, high-resource languages have the highest scores in supervised and crosslingual settings. Overall, we report relatively higher correlation scores which vary between 0.40 and 0.88.
5 Conclusion
We presented SemRel, a new collection of semantic textual relatedness datasets in 13 languages with the majority predominantly spoken in Africa and Asia and considered low-resource. The sentence pairs contained in the datasets are annotated by native speakers and are associated with fine-grained relatedness scores. We reported the details related to the data curation and emphasised the challenges faced when dealing with low-resource languages.
We publicly release the datasets as well as other resources, such as the annotation guidelines and full labels for the research community interested in semantic relatedness, low-resource languages, and disagreements.
6 Limitations
We acknowledge that there is no formal definition of what constitutes semantic relatedness. Hence, the annotations may be subjective. To mitigate the issue we share our guidelines and annotated instances so researchers in the community can expand on our work, replicate, and study the disagreements in our data. We are also aware of the limited number of data sources and data variety in some low-resource languages involved. We do not claim that the datasets released represent all variations of these languages but they remain a good starting point as they were carefully picked, labelled, and processed by native speakers.
Although our collection is comprised of multiple datasets, the size of the data is limited, thus it cannot be the only source used for tasks that require a large amount of data such as language identification.
7 Ethical Considerations
All the annotators involved in this study were either volunteers or were paid more than the minimum wage per hour and any demographic information reported in the Appendix was shared with consent. The data that was further annotated was publicly available and is cited in our paper.
Similarly to Abdalla et al. (2023), we acknowledge all the possible socio-cultural biases that can come with our data, due to the data sources or the annotation process. When building our datasets, we did avoid instances with inappropriate or offensive utterances but we might have missed some. Our goal was to identify common perceptions of semantic relatedness by native speakers and our labels are not meant to be standardised for any given language. Note that we build datasets for low-resource languages but we do not claim in any way that these are fully representative of their usage.
Acknowledgements
We thank our annotators for labelling the data and for the insightful comments as well as Zara Siddique for providing additional insights.
Thanks to Dimosthenis Antypas, Joanne Boisson, Hsuvas Borkakoty for the helpful feedback.
Appendix
Appendix A Annotation
Lang. Datasets (%)
afr Oscar (100%).
amh News data (100%).
arb Ted Talk subtitles on science, art, and society (96%), Economy news data (4%).
arq Conversational data (89%), Youtube comments (11%).
ary News data (100%).
eng Wikipedia (29%), ParaNMT (17%), Formality (17%), SNLI (8%), Goodreads (22%), STS (7%).
esp MuchoChine (MC) (9%), Spanish QC (4%), PAWS-X Paraphrasing dataset (16%), NLI-es (11%),
SICK-es (16%), STS (5%), SP (19.5%), XL-Sum (19.5%).
hau News data (100%)
hin News data (100%).
ind News data (82%), Wikipedia/ROOTS (18%).
kin News data (100%).
mar News data (100%).
tel News data (100%).
Table 6: The percentage of instances collected from different sources (datasets).
A.1 Pilot data annotation
To assess the different pairing techniques and the potential annotation challenges, we run a pilot annotation task on 20 to 100 pairs of sentences for each language before proceeding with larger annotation batches. This helped us assess the difficulties related to the annotation task and the choices to be made for the final data processing step. For instance, if highly related and unrelated pairs were occurring too often in the tuples, we reduced the percentages of both highly related and unrelated pairs by changing or calibrating the data sources if possible, prioritising other pairing techniques, or including an extra preprocessing step (e.g., paraphrase detection).
A.2 Data Pre-processing Tools
We used NLTK tools for parsing Afrikaans and Indonesian in addition to manual verification. For instance, as for Indonesian, the NLTK sentence parses generated many errors due to common Indonesian abbreviations that involve ’.’, which the sentence parser mistakenly detects as the end of a sentence, we added new abbreviations for parsing (’ir.’, ’kh.’, ’h.’, ’drs.’, ’drg.’, ’rm.’, ’bp.’, ’bpk.’, ’tgl.’, ’no.’, ’jl.’, and ’jln.’)
A.3 Information about the Annotators
We report on the demographic information of the volunteers who agreed to share them.
Afrikaans
Paid native speakers.
Amharic
Paid Amharic native speakers, 3 women and 5 men from different social, cultural, and ethnic backgrounds (Amhara, Guragie, Wolyta, Sidama, and Oromo).
Modern Standard Arabic and Algerian Arabic
Native speakers, 2 men, 2 women, university degree holders, ages vary between 23 to 56, paid above the minimum wage.
Moroccan Arabic
Volunteer native speakers, 3 women, 1 man, university degree holders, volunteers.
English and Spanish
Amazon Mechanical Turkers with high approval rates (98% for English) paid above the US minimum wage.
Hausa
Paid native speakers, 3 women, 1 man, age: 28 to 30, bachelor’s degree holders.
Hindi and Marathi
Paid native speakers.
Telugu
Volunteer native speakers.
A.4 Annotation Guidelines
You will be given four sentence pairs (i.e., 4 pairs of the form [sentence A, sentence B]). Your task is to judge the relatedness of each pair (sentence A and sentence B) and tell us:
  • the sentence pair that is the MOST related (i.e., sentence A is closest in meaning to sentence B).
  • the sentence pair that is the LEAST related (i.e., sentence A is farthest in meaning to sentence B).
Sentence pairs can be related in many ways. I.e., sentence A and sentence B can be related in different ways. The first pair of sentences in Table 7 are more related than the second one.
MOST Related Pair S1: The boy enjoyed reading under the lemon tree
S2: There is a lemon tree next to the house
LEAST related Pair S1: The boy enjoyed reading under the lemon tree
S2: The boy was an excellent football player
Table 7: Example in the Guidelines. Examples of two pairs of sentences with different degrees of relatedness from Abdalla et al. (2023).
Often, sentence pairs that are more specific in what they share tend to be more related than sentence pairs that are only loosely about the same topic.
If a sentence has more than one interpretation, consider that meaning which is closest to the meaning of the other sentence in the pair. If both sentences have multiple meanings, then consider those meanings that are closest to each other.
Pair 1 S1: The boy enjoyed reading under the lemon tree
S2: I have a green hat
Pair 2 S1: The boy enjoyed reading under the lemon tree
S2: She was an excellent football player
Table 8: Example in the Guidelines. Examples of two pairs of sentences that have similar degrees of relatedness where one can choose randomly the most vs. least related pairs (i.e., either Pair 1 or Pair 2).
If in the given set of four pairs, two (or more) sentence pairs are equally related to each other and they are also the most related pairs, then select either one of them as the most related (i.e., randomly). Similarly, if two (or more) equally related pairs are also the least related pairs, then select either one of them as the least related. (See Table 2.)
You cannot select the same sentence pair for both categories.
Try not to overthink the answer. Let your instinct guide you.
A.5 Notes
Sentence pairs can be related in many ways. Consider the entire meaning of the sentences before selecting the most related. The sentences included in this task may contain foul language, though we have attempted to limit this.
A.6 Examples (Q1)
Which of the four sentence pairs in Table 9 is MOST RELATED? Which pair is LEAST RELATED?
Pair 1 S1: My personal favorites from Narnia were the conversations between Aslan and Bree.
S2: This marks my progress through the Chronicles, picked up after reading The Narnia Code and Planet Narnia.
Pair 2 S1: why won’t she ask me out?
S2: and after all that you wont have to worry about getting a girl to like you.
Pair 3 S1: A group of people are sitting on the grass outside of a rustic building.
S2: Group sitting on a grassy hill resting.
Pair 4 S1: If you change me back, I will feed each one of your snakes a large mouse!
S2: Offer people who join cash and coupons.
Table 9: Q1 Example in the Guidelines.
A.6.1 A1
The most related pair is Pair 3 because both sentences are talking about a group sitting/resting in grass.
The least related is Pair 4 because Pair 4 sentences are completely unrelated, whereas the other pairs have some relatedness.
A.6.2 Note (A1)
Pair 1 sentences are somewhat related, as they talk about Narnia/characters in that world (Aslan and Bree are characters in Narnia). However, the content of this sentence pair is not as related as Pair 3.
Pair 2 sentences are both talking about romantic relationships.
A.7 Examples (Q2)
Which of the four sentence pairs in Table 10 is MOST RELATED? Which pair is LEAST RELATED?
Pair 1 S1: That and a kids meal.
S2: My two kids, ages 5 and 3!
Pair 2 S1: The spines , which may be up to 50 mm long , are modified hairs , mostly made of keratin .
S2: The simplest shape is the long opening with a pointed arch known in England as the lancet .
Pair 3 S1: A woman wearing a white shirt and a red headband is sitting outside.
S2: Two women stand outside a library.
Pair 4 S1: Ayodhya ,capital of King Rama is mentioned on the banks of Sarayu river .
S2: Ramayana mentions that city of Ayodhya was situated on the bank of Sarayu river .
Table 10: Q2 Example in the Guidelines.
A.7.1 A2
The most related pair is Pair 4. Both sentences are talking about the same city and mention that it is on the bank of river Sarayu. The least related pair is Pair 2 because the sentences are completely unrelated.
A.7.2 Note (A2)
Pair 3 sentences both refer to at least one woman outside.
Pair 1 sentences refer to kids or kid-related things (making them slightly close in meaning).
A.8 Examples (Q3)
Which of the four sentence pairs in Table 11 is MOST RELATED? Which pair is LEAST RELATED?
Pair 1 S1: IBM has not shifted its focus from mainframes to compete with Windows
S2: In 3 years, IBM has not been interested in the PC.
Pair 2 S1: I wanted to see the scene where Quinn told the brotherhood he was in love with Blay.
S2: I also would have liked to see the scene where Qhuinn asks Blay’s dad for permission to propose to Blay.
Pair 3 S1: Jeremy desperately needs a stable home.
S2: Furnishings were an angle bed, a stool, and a chamber pot on the dirt floor.
Pair 4 S1: That’s difficult. They’re both great
S2: that’s really hard they are both great!
Table 11: Q3 Example in the Guidelines.
A.8.1 A3
The most related pair is Pair 4. Both sentences are paraphrases of each other. (Pair 1 and Pair 2 are quite related but not as exact paraphrases as Pair 4.)
The least related pair is Pair 3. Pair 3 sentences are somewhat related as they talk about house furnishings. However, they are still less related than all the other pairs.
A.8.2 Note (A3)
Pair 1 sentences both refer to IBM and their business strategy. We consider this to be more related than Pair 3 because it’s more specific in the details they share.
Pair 2 sentences talk about the same characters and their romantic situation.
Appendix B Pre-trained models used
We list down the various pre-trained HuggingFace models used in our experiments:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
Language Base Finetuned
afr 0.76 0.79
amh 0.79 0.85
arb 0.55 0.62
arq 0.40 0.60
ary 0.38 0.77
eng 0.82 0.83
esp 0.65 0.70
hau 0.48 0.69
hin 0.71 0.77
ind 0.53 0.50
kin 0.45 0.72
mar 0.82 0.88
tel 0.80 0.82
Table 12: Spearman correlation scores on LaBSE models with and without further fine-tuning on our training data (Base and fine-tuned, respectively).
Language mBERT XLMR
afr 0.77 0.76
amh 0.12 0.69
arb 0.40 0.42
arq 0.28 0.32
ary 0.53 0.50
eng 0.71 0.74
esp 0.67 0.68
hau 0.32 0.31
hin 0.64 0.63
ind 0.54 0.54
kin 0.25 0.30
mar 0.78 0.75
tel 0.77 0.78
Table 13: Spearman correlation of the BERTScore Zhang* et al. (2020) with mBERT and XLMR on the different languages.

### v5

@@
 (2017a).
-Similarly, we selected Spanish sentence pairs from semantic similarity datasets Agirre et al.
+Similarly, we select pairs of Spanish sentences from semantic similarity datasets such as STS Agirre et al.
 (2014, 2015); Cer et al.
@@
 (2021) and NLI-es Araujo et al.
+(2022), and paraphrasing datasets such as PAWS-X Yang et al.
 (2019).
-Additionally, we sampled contiguous sentences from the summarisation dataset XL-Sum Hasan et al.
-(2021), DiscoEval Spanish Araujo et al.
-(2022), and questions of different types from Spanish QC Á.
+We also sampled contiguous sentences from XL-Sum Hasan et al.
+(2021) and BSO DiscoEval Spanish Araujo et al.
+(2022), and we included questions of different types from Spanish QC Á.
 García Cumbreras et al.
@@
 (2020), respectively.
-Sentences shorter than five words and longer than 20 were excluded and pairs were created using lexical overlap.
+Sentences shorter than five words and longer than 20 were excluded, and pairs were created using lexical overlap.
 Additionally, for Amharic, we excluded sentences with mixed languages to avoid confusing the annotators.
@@
 (2022); Setya and Mahendra (2018) and the IndoSum Kurniawan and Louvan (2018) datasets.
-IndoSum is a human-written summarization dataset consisting of pairs of news articles with their abstractive summaries.
-We parsed both corpora at a sentence-level and only selected sentences that are composed of five to fifteen words.
+IndoSum is a human-written summarization dataset consisting of pairs of news articles with abstractive summaries.
+We parsed both corpora at a sentence level and only selected sentences that were composed of five to fifteen words.
 Hindi, Marathi, and Telugu
@@
 We inspected annotators with large disagreements to ensure the annotation procedure was correctly followed (i.e., their annotations made sense for native speakers).
-Very strong disagreements would serve as a red flag of poor data quality resulting in a more thorough review of the annotator quality.
-Hence, as a sanity check, we looked to see that sentences with high relatedness scores were more semantically related than those with low relatedness scores.
+Very strong disagreements would serve as a red flag of poor data quality resulting in a more thorough review of the annotation quality.
+Hence, as a sanity check, we examined whether sentences with high relatedness scores were more semantically related than those with low relatedness scores.
 The specific procedure for ensuring data quality depends on the annotation procedure of the team (e.g., those using AMT vs those who did not).
@@
 Manual Spotchecks
-Finally, a team of native speakers manually spotchecked the scores to make sure that the relatedness scores made sense and to supplement the quantitative evaluation based on SHR.
+Finally, a team of native speakers manually spot-checked the scores to make sure that the relatedness scores made sense and to supplement the quantitative evaluation based on SHR.
 4 Experiments
@@
 eng Wikipedia (29%), ParaNMT (17%), Formality (17%), SNLI (8%), Goodreads (22%), STS (7%).
-esp MuchoChine (MC) (9%), Spanish QC (4%), PAWS-X Paraphrasing dataset (16%), NLI-es (11%),
-SICK-es (16%), STS (5%), SP (19.5%), XL-Sum (19.5%).
+esp MuchoChine (MC) (7%), Spanish QC (4%), PAWS-X (19.5%), NLI-es (12%), SICK-es (17%),
+STS (5%), BSO (17.5%), XL-Sum (18%).
 hau News data (100%)